# DRWKV: Focusing on Object Edges for Low-Light Image Enhancement

### v2

@@ Abstract
Low-light image enhancement remains a challenging task, particularly in preserving object edge continuity and fine structural details under extreme illumination degradation. In this paper, we propose a novel model, DRWKV (Detailed Receptance Weighted Key Value), which integrates our proposed Global Edge Retinex (GER) theory, enabling effective decoupling of illumination and edge structures for enhanced edge fidelity. Secondly, we introduce Evolving WKV Attention, a spiral-scanning mechanism that captures spatial edge continuity and models irregular structures more effectively. Thirdly, we design the Bilateral Spectrum Aligner (Bi-SAB) and a tailored MS²-Loss to jointly align luminance and chrominance features, improving visual naturalness and mitigating artifacts. Extensive experiments on five LLIE benchmarks demonstrate that DRWKV achieves leading performance in PSNR, SSIM, and NIQE while maintaining low computational complexity. Furthermore, DRWKV enhances downstream performance in low-light multi-object tracking tasks, validating its generalization capabilities.The code are available at: https://github.com/JackBaixue/DRWKV
Index Terms:
Low-light Image Enhancement, Receptance Weighted Key Value, Global Edge Retinex theory
I Introduction
Low-Light Image Enhancement (LLIE) technology is regarded as an active method to brighten low-light images at low cost [1, 2]. Recently, the academic community has primarily explored prominent methods such as global attention modeling [3, 4, 5], local feature enhancement [6, 7, 8], and selective state updating [9, 10]. Based on the robust Retinex theory [11], the composition of a low-light image $I$ can be interpreted as:
@@ I Introduction
where $R$ denotes scene reflectance, $S$ represents the illumination component, and $N$ stands for intrinsic noise.
@@
When constructing a low-light scene, for the image, in addition to the increase in noise $N$, the illumination component $L$ is decomposed into global and local illumination. The global spectrum and contour light field will dynamically change under light reflection. In the context of low-light image quality improvement, focusing solely on global illumination will amplify background noise [12] and cause edge distortion [13], while focusing only on contours leads to abrupt transitions and globally underexposed results. Therefore, under the condition that global exposure is controllable, the essence of low-light image enhancement lies in the refinement of object edges. As shown in Figure 1(c), ZeroDCE [8] suffers from edge distortion, and RetinexMamba [14] exhibits detail blurring, indicating that current methods have limitations in handling extremely low-light environments.
To address these issues, we propose Detailed Receptance Weighted Key Value (DRWKV) model, which optimizes the Bi-WKV Attention of VRWKV (in ICLR 2025) [15]—a classic model in the visual RWKV domain. We introduce Evolving WKV Attention to enhance the extraction of fine structures in object edges. Compared with Bi-WKV Attention, the spatial topological awareness of Evolving WKV Attention can maximize the capture of continuous spaces of object edges, thereby strengthening the modeling of irregular spatial structures. Following this, we design an auxiliary technique—Bilateral Spectrum Aligner (Bi-SAB). By aligning brightness and color spectral features, SpAligner resolves issues of over-exposure and color distortion caused by separate processing. Coupled with MS²-Loss, SpAligner renders the illumination and colors of enhanced images more natural. We conducted extensive image-enhancement and low-light object-tracking experiments on five benchmark datasets. As illustrated in Figure 1(a) and 1(b), our method achieves leading PSNR, SSIM, and NIQE scores, demonstrating significant superiority in edge continuity and detail fidelity.
@@
  • We introduce Bilateral Spectrum Aligner (Bi-SAB), which enriches the color domain of brightened low-light images by coordinating spectral features and coupling with our proposed MS²-Loss.
II Related Work
II-A Object Edge Perception and Extraction
Edge perception and extraction techniques are built upon traditional methods [16, 17, 18, 19] and deep learning approaches [20, 8]. They achieve fine-grained restoration of object edges in low-light scenarios by fusing decoupled edge features with contextual information.
Different from traditional CNN-based methods such as Retinex-Net [21] and LLNet [22] that rely on local convolutions to extract edge features, emerging edge-aware approaches introduce topological modeling [23, 24, 25]. These methods treat edges as ”spatial sequences” and capture cross-regional features through state accumulation. This design offers a core advantage in enhancing weak edge signals by achieving balanced transitions and local optimization through the adjustment of edge gradient consistency constraints and dynamic thresholds.
Figure 2: The overview of the DRWKV model. (a) DRWKV processes low-light images through two steps: (i) Light Preprocessing and (ii) Deep Detail Mining. The positions of Block1 and Block2 correspond to the layout settings of the two Retinex theories. (b) The block structure of Evolving Scanning RWKV (ES-RWKV), which embeds two core mixing modules: Evolving Scanning Spatial Mix and Channel Mix. (c) The block structure of the Bilateral Spectrum Aligner Block (Bi-SAB).
II-B Low-Light Object Edge Restoration and Enhancement
@@
Techniques such as EDTER [26] and Retinex-RAWMamba [27] explore diverse architectures by leveraging self-attention to capture cross-regional dependencies and linear scanning to optimize temporal accumulation.
Robust edge extraction is achieved through frequency-domain feature filtering [28], multi-modal edge complementarity [29, 30], and adversarial learning-based edge constraints [14], further precisely enhancing edge features.
@@ III Methodology
(1) Neural Architecture. VRWKV adopts a single recursive structure, which mismatches the hierarchical features of edges. This causes aliasing of shallow and deep features (manifested as edge blurring), fails to decouple illumination variations from edge representation, and leads to frequent misjudgment between noise and edges in bright/dark regions under low light.
@@
Based on the results of our reasoning, we propose two hypotheses: (1) Hierarchical detail enhancement depends on a gradient-guided mechanism with illumination invariance. (2) Modeling the continuity of low-light edges requires a temporal encoding scheme compatible with the spatial topology of edges.
DRWKV Model: Different from existing mainstream methods, we adopt a layered optimization strategy, which prioritizes combining the gray-world assumption with the three-branch channel consisting of noise N, artifact S, and illumination L to implement (III-A) Light Preprocessing, thereby obtaining the brightness-restored image $\hat{I}$. This follows a strategy similar to that of Retinexformer [31]. The RWKV structure comprises a set formed by four groups of ES-RWKV and sampling blocks. For the encoder, (III-B) Evolving Scanning is applied for the first time, combined with wavelet transform downsampling to accomplish edge gradient feature extraction. The decoder integrates the sampling information $F$ and $F^{\prime}$ via the (III-C) Bilateral Spectrum Aligner Block (Bi-SAB), enabling progressive enhancement from microscopic textures to macroscopic contours. The new design accounts for nonlinear edge topology and signal-to-noise ratio constraints, balancing the depth of low-light detail mining and computational cost. The overall structure of our DRWKV model is provided in Figure 2.
III-A Light Preprocessing
The Light Preprocessing structure employs a dual-path parallel guidance approach. Technically, we recalculate the RGB components using the gray-world assumption $L_{ill}$:
$L_{\text{ill}}=\begin{bmatrix}R^{\prime}\ G^{\prime}\ B^{\prime}\end{bmatrix}=\text{diag}\left(\frac{\frac{1}{3}\sum_{i=1}^{N}\left|c_{i}\right|1}{\sum{i=1}^{N}c_{i}}\right)\cdot c$ (2)
Here, $c_{i}=[R_{i},G_{i},B_{i}]^{T}$ restructures the three-channel vector of a pixel. The detailed procedure involves converting the vector into a diagonal matrix via $\text{diag}(\cdot)$ and then computing the inner product with the original channel vector $c$. By combining $L_{ill}$ with the artifact component $S$, we explicitly define the interference boundaries to be avoided, enabling the Global Edge Retinex (GER) to enhance brightness without amplifying artifacts, thus achieving the first path of clean baseline enhancement.
@@ III-A Light Preprocessing
$\hat{I}=I\times\hat{R}=I\times(I-N)/L$ (3)
In the second path, we process different information components by computing the product of the reconstructed reflectance component $\hat{R}$ and the input low-light image $I$, resulting in a low-light enhanced image $\hat{I}$ that retains the original detail information.
In summary, Light Preprocessing balances brightness, preserves details, and ensures the baseline quality of the low-light enhanced results for the subsequent RWKV stage.
III-B Deep Detail Mining
@@
The robust Retinex theory suffers from limitations: lack of hierarchical edge representation and strong coupling between illumination and edges. Hypothesis(1) posits that under low-light conditions, the key to detail optimization lies in decoupling components under an illumination-invariant feature basis, combined with gradient information for differential enhancement of shallow and deep details. We propose the Global Edge Retinex (GER) Theory to substantiate this, with its core formulation as:
$I=(R+\alpha\cdot E)\odot L+\beta\cdot N+\gamma\cdot S$ (4)
Compared with traditional Retinex models, GER’s improvements lie in three aspects: introducing a new edge feature term $\alpha E$ to model hierarchical details; incorporating an artifact term $\gamma S$ to suppress reconstruction noise; and redefining $N$ as spatially heterogeneous noise.
@@
$R=\frac{I-N}{L+\epsilon}\in[0,1],$ (5)
Here, $R$ strips off spatially heterogeneous noise via INI-N and neutralizes illumination interference through $L+\epsilon$, providing a ”clean” benchmark for detail enhancement.
To address shallow detail blurring, GER introduces the edge feature $E$ to construct the enhanced reflectance map:
$R_{\text{enh}}=R+\alpha\cdot E,$ (6)
With $R$ as the anchor, $E$ guides differential enhancement of shallow and deep details, mitigating feature aliasing caused by the single recursion in VRWKV, while incorporating gradient information to strengthen edge continuity.
The illumination component $L$ is estimated via the gray-world assumption in the Light Preprocessing step: $L=\text{Illumination}(I)\in(0,2)$. To resolve edge misjudgment in bright and dark regions under low light, GER sets the artifact residual suppression term $S_{\text{supp}}=\gamma\cdot S$.
Finally, the enhanced image is obtained through synergistic estimation of all components:
$\hat{I}=R_{\text{enh}}\times L+\beta\cdot N+S_{\text{supp}}.$ (7)
@@
Evolving Scanning (ES) technology addresses Hypothesis(2) by transforming the geometric continuity of edge manifolds into temporal continuity capturable by the VRWKV propagation mechanism. Its mathematical foundation lies in the Archimedean spiral equation:
@@
Here, $r(\theta)=a+b\theta$ governs the initial radius ($a$) and expansion rate ($b$) of the spiral. The path expands uniformly from the starting point, naturally adapting to the edge characteristics of spreading from local to global regions. Notably, we extend a single spiral to a four-directional spiral system, which initiates from the four corners of the feature map and supports both clockwise and counterclockwise directions, achieving targeted adaptation to low-light edge issues.
To ensure that the spiral path preserves the geometric adjacency of edge points, strict topology-preserving constraints are introduced. The constraint rule is as follows: for any two points $e_{i},e_{j}$ in the edge point set $E$, if their Riemannian geometric distance satisfies $\|e_{i}-e_{j}\|_{\text{geo}}<\delta$ (where $\delta$ is the geometric proximity threshold), their temporal positions $t_{i},t_{j}$ in the one-dimensional sequence generated by the spiral path must satisfy $|t_{i}-t_{j}|<\tau$ (where $\tau$ is the temporal proximity threshold).
At the network level, ES-RWKV is a fusion of ES technology and the RWKV structure. It bridges the gap between the Riemannian manifold nature of low-light edges and the temporal modeling capability of VRWKV, effectively enhancing the modeling performance of edge continuity.
III-C Bilateral Spectrum Aligner
@@
5:/*Computer Cross Attention (CA) feature representations.*/
@@
7:/*Computer Feature Difference Adjustment (FDA) feature representations.*/
@@
13:/*Computer Spectral Alignment Enhancer (SAE) feature representations.*/
@@
15:/*Computer Bi-SAB feature representations.*/
@@
III-D Loss Function
TABLE I: Experiment results on the LSRW-Huawei, LOLv2, and SDSD test benchmarks. The background color gradients from dark to light, with red indicating the optimal result, orange denoting the sub-optimal result, and yellow following in order.
Method GFLOPs LSRW-Huawei LOLv2_Real LOLv2_Syn SDSD_indoor SDSD_outdoor

### v3

@@ Abstract
Low-Light Image Enhancement (LLIE) remains a challenging task, particularly in preserving object edge continuity and fine structural details under extreme illumination degradation. In this paper, we propose a novel model, DRWKV (Detailed Receptance Weighted Key Value), which integrates our proposed Global Edge Retinex (GER) theory, enabling effective decoupling of illumination and edge structures for enhanced edge fidelity. Secondly, we introduce Evolving WKV Attention, a spiral-scanning mechanism that captures spatial edge continuity and models irregular structures more effectively. Thirdly, we design the Bilateral Spectrum Aligner Block (Bi-SAB) and a tailored MS2-Loss to jointly align luminance and chrominance features, improving visual naturalness and mitigating artifacts. Extensive experiments on five LLIE benchmarks demonstrate that DRWKV achieves leading performance in PSNR, SSIM, and NIQE while maintaining low computational complexity. Furthermore, DRWKV enhances downstream performance in low-light multi-object tracking tasks, validating its generalization capabilities. The code are available at: https://github.com/JackBaixue/DRWKV
1 Introduction
Low-Light Image Enhancement (LLIE) technology is regarded as an active method to brighten low-light images at low cost [nguyen2024diffusion, Tan_2024_ACCV]. Recently, the academic community has primarily explored prominent methods such as global attention modeling [xu2025upt, wang2024low, shang2024holistic], local feature enhancement [zhang2024cnn, guo2016lime, guo2020zero], and selective state updating [zou2024wave, qu2024double]. Based on the robust Retinex theory [8304597], the composition of a low-light image $I$ can be interpreted as:
@@ 1 Introduction
where $R$ denotes scene reflectance, $L$ represents the illumination component, and $N$ stands for intrinsic noise.
@@
When constructing a low-light scene, for the image, in addition to the increase in noise $N$, the illumination component $L$ is decomposed into global and local illumination. The global spectrum and contour light field will dynamically change under light reflection. In the context of low-light image quality improvement, focusing solely on global illumination will amplify background noise [shi2024zero] and cause edge distortion [wang2024multimodal], while focusing only on contours in isolation leads to rigid edge transitions. Therefore, under the condition that global exposure is controllable, the essence of low-light image enhancement lies in the refinement of object edges. As shown in Figure 1(c), ZeroDCE [guo2020zero] suffers from edge distortion, and RetinexMamba [bai2024retinexmamba] exhibits detail blurring, indicating that current methods have limitations in handling extremely low-light environments.
To address these issues, we propose Detailed Receptance Weighted Key Value (DRWKV) model, which optimizes the Bi-WKV Attention of VRWKV (in ICLR 2025) [duan2024vision]—a classic model in the visual RWKV domain. We introduce the Evolving WKV Attention, a novel attention mechanism with superior spatial topological awareness, which can maximize the capture of continuous spatial information of object edges, thereby strengthening the modeling effect of irregular spatial structures. Following this, we design an auxiliary technique—Bilateral Spectrum Aligner Block (Bi-SAB). By aligning brightness and color spectral features, Bi-SAB resolves issues of over-exposure and color distortion caused by separate processing. Coupled with MS²-Loss, Bi-SAB renders the illumination and colors of enhanced images more natural. We conducted extensive image-enhancement and low-light object-tracking experiments on five benchmark datasets. As illustrated in Figure 1(a) and 1(b), our method achieves leading PSNR, SSIM, and NIQE scores, demonstrating significant superiority in edge continuity and detail fidelity.
@@
  • We introduce Bilateral Spectrum Aligner Block (Bi-SAB), which enriches the color domain of brightened low-light images by coordinating spectral features and coupling with our proposed MS²-Loss.
2 Related Work
2.1 Object Edge Perception and Extraction
Edge perception and extraction techniques are built upon traditional methods [chen2024highly, lu2022low, 9328179, lim2020dslr] and deep learning approaches [wang2022eanet, guo2020zero]. They achieve fine-grained restoration of object edges in low-light scenarios by fusing decoupled edge features with contextual information.
Different from traditional CNN-based methods such as Retinex-Net [wei2018deep] and LLNet [lore2017llnet] that rely on local convolutions to extract edge features, emerging edge-aware approaches introduce topological modeling [zhao2024dyedgegat, li2023edge, tang2025spiralmamba]. These methods treat edges as “spatial sequences” and capture cross-regional features through state accumulation. This design offers a core advantage in enhancing weak edge signals by achieving balanced transitions and local optimization through the adjustment of edge gradient consistency constraints and dynamic thresholds.
Figure 2: The overview of the DRWKV model. (a) DRWKV processes low-light images through two steps: (i) Light Preprocessing and (ii) Deep Detail Mining. The positions of Block1 and Block2 correspond to the layout settings of the two Retinex theories. (b) The block structure of Evolving Scanning RWKV, which embeds two core mixing modules: Evolving Scanning Spatial Mix and Channel Mix. (c) The block structure of the Bilateral Spectrum Aligner Block.
2.2 Low-Light Object Edge Restoration and Enhancement
@@
Techniques such as EDTER [pu2022edter] and Retinex-RAWMamba [chen2024retinex] explore diverse architectures by leveraging self-attention to capture cross-regional dependencies and linear scanning to optimize temporal accumulation.
Robust edge extraction is achieved through frequency-domain feature filtering [guan2021dual], multi-modal edge complementarity [qian2021improved, liu2024optical], and adversarial learning-based edge constraints [bai2024retinexmamba], further precisely enhancing edge features.
@@ 3 Methodology
(1) Neural Architecture. VRWKV adopts a single recursive structure, which mismatches the hierarchical features of edges. This causes aliasing of shallow and deep features, manifested as the failure to decouple illumination variations from edge representation, and leads to frequent misjudgment between noise and edges in low-light environments.
@@
Based on the results of our reasoning, we propose two hypotheses: (1) Hierarchical detail enhancement depends on a gradient-guided mechanism with illumination invariance. (2) Modeling the continuity of low-light edges requires a spatiotemporal alignment modeling scheme compatible with the topological properties of edges.
DRWKV Model: Different from current mainstream methods, model adopts a layered optimization strategy, which prioritizes combining the gray-world assumption with a three-branch channel consisting of noise $N$, artifact $S$, and illumination $L$ to implement (3.1) Light Preprocessing (LP), thereby obtaining the brightness-restored image $\hat{I}$ [cai2023retinexformer]. For the Deep Detail Mining (DDM) process, the RWKV structure comprises a set formed by four groups of ES-RWKV and sampling blocks. For the encoder, (3.2) Evolving Scanning mechanism is applied for the first time, combined with wavelet transform downsampling to accomplish edge gradient feature extraction. The decoder integrates the sampling information $F$ and $F^{\prime}$ via the (3.3) Bilateral Spectrum Aligner Block (Bi-SAB), enabling progressive enhancement from microscopic textures to macroscopic contours. The overall structure of our DRWKV model is provided in Figure 2.
3.1 Light Preprocessing
The Light Preprocessing (LP) structure employs a dual-path parallel guidance approach, balancing brightness, preserving details, and ensuring the baseline quality of LLIE results for the subsequent Deep Detail Mining (DDM) process. Technically, we recalculate the RGB components using the gray-world assumption $L_{ill}$:
$L_{\text{ill}}=\begin{bmatrix}R^{\prime}\ G^{\prime}\ B^{\prime}\end{bmatrix}=\text{diag}\left(\frac{\frac{1}{3}\sum_{i=1}^{N}\left|c_{i}\right|}{\sum_{i=1}^{N}c_{i}}\right)\cdot c,$ (2)
where $c_{i}=[R_{i},G_{i},B_{i}]^{T}$ restructures the three-channel vector of a pixel. The detailed procedure involves converting the vector into a diagonal matrix via $\text{diag}(\cdot)$ and then computing the inner product with the original channel vector $c$. By combining $L_{ill}$ with the artifact component $S$, we explicitly define the interference boundaries to be avoided, enabling the Global Edge Retinex (GER) to enhance brightness without amplifying artifacts, thus achieving the first path of clean baseline enhancement.
@@ 3.1 Light Preprocessing
$\hat{I}=I\times\hat{R}=I\times(I-N)/L.$ (3)
In the second path, we process different information components by computing the product of the reconstructed reflectance component $\hat{R}$ and the input low-light image $I$, resulting in a low-light enhanced image $\hat{I}$ that retains the original detail information. lts for the subsequent RWKV stage.
3.2 Deep Detail Mining
@@
The traditional Robust Retinex theory has limitations in component decoupling and deep/shallow detail processing. Based on the hierarchical detail enhancement approach in Hypothesis(1), we propose the Global Edge Retinex (GER) theory, and the core formula is as follows:
$I=(R+\alpha\cdot E)\odot L_{ill}+\beta\cdot N+\gamma\cdot S.$ (4)
The theory takes the explicit edge feature $E$ (the formula is provided in Appendix 3.3), the spatially heterogeneous noise $N$ (obtained after segmenting traditional noise), the gray-world assumption $L_{ill}$, the artifact $S$, and the learnable parameters $\alpha$, $\beta$, $\gamma$ (which control feature expression) as the anchors for addressing low-light edge detail issues.
@@
$R=\frac{I-N}{L_{ill}+\epsilon}\in[0,1],$ (5)
where $R$ strips off spatially heterogeneous noise via INI-N and neutralizes illumination interference through $L_{ill}+\epsilon$, providing a ”clean” benchmark for detail enhancement.
From the model perspective, the edge feature $E$ (processed by spiral gradient scanning) is used to construct the enhanced reflectance map:
$R_{\text{enh}}=R+\alpha\cdot E.$ (6)
The parameter $\alpha$ is used to control the intensity of differential enhancement of deep/shallow details guided by $E$, ensuring clear feature hierarchy. Meanwhile, gradient information is incorporated to enhance edge continuity.
Finally, an enhanced image with edge fidelity is obtained through the collaborative balance of all components:
$\hat{I}=R_{\text{enh}}\times L_{ill}+\beta\cdot N+\gamma\cdot S.$ (7)
@@
Evolving Scanning (ES) mechanism addresses Hypothesis(2) by transforming the geometric continuity of edge manifolds into temporal continuity capturable by the VRWKV propagation mechanism. Its mathematical foundation lies in the Archimedean spiral equation:
@@
where $r(\theta)=a+b\theta$ governs the initial radius ($a$) and expansion rate ($b$) of the spiral. The path expands uniformly from the starting point, naturally adapting to the edge characteristics of spreading from local to global regions. In terms of the scanning mechanism, we extend a single spiral to a four-directional rectangular spiral system. Let $P$ denote the spiral path spreading outward from the center, and let $P_{i}=(x_{i},y_{i})$ represent the coordinates of the $i$-th step (where $i\in\mathbb{N}$). The coordinates satisfy:
$\begin{split}P_{i}&=(x_{0},y_{0})+\left(r_{0}+\alpha\lfloor i/4\rfloor\right)\cdot\delta_{imod4},\\ &\quad\text{where }\delta\in\{(0,-1),(1,0),(0,1),(-1,0)\}.\end{split}$ (9)
Here, $(x_{0},y_{0})$ is the spiral center, $r_{0}>0$ is the initial radius, and $\alpha>0$ is the radial pitch. The direction vector $\delta_{imod4}$ cycles incrementally in the four directions of up, right, down, and left.
In terms of implementation, the spiral path is encoded into several token streams, which are input into the Bi-WKV attention mechanism for iterative modeling. By applying the proposed Evolving Scanning mechanism to the RWKV structure, we ultimately bridge the gap between the Riemannian manifold properties of low-illumination edges and the temporal modeling capability of RWKV. This significantly improves edge continuity and structural fidelity under low-light conditions (the formula derivation process is provided in Appendix 2):
$ES\text{-}RWKV(\cdot)=\cdot+\mathcal{F}_{\text{spatial}}(\cdot)+\mathcal{F}_{\text{channel}}(F_{\text{ESSM}}),$ (10)
where $\cdot$ denotes the input feature, $\mathcal{F}_{\text{spatial}}(\cdot)$ represents the spatial dependency modeling implemented via spiral scanning and the ES-RWKV mechanism, and $\mathcal{F}_{\text{channel}}(\cdot)$ denotes the channel correlation modeling based on Q-Shift and nonlinear mapping.
3.3 Bilateral Spectrum Aligner Block
@@
5:/*Compute Cross Attention (CA) feature representations.*/
@@
7:/*Compute Feature Difference Adjustment (FDA) feature representations.*/
@@
13:/*Compute Spectral Alignment Enhancer (SAE) feature representations.*/
@@
15:/*Compute Bi-SAB feature representations.*/
@@
3.4 Loss Function
Table 1: Experiment results on the LSRW-Huawei, LOLv2, and SDSD test benchmarks. The background color gradients from dark to light, with red indicating the optimal result, orange denoting the sub-optimal result, and yellow following in order.
Method Params GFLOPs LSRW-Huawei LOLv2_Real LOLv2_Syn SDSD_indoor SDSD_outdoor